OpenStudy (christos):

Solve the limit limit(x-->+infinity) sqrt((x-3)/x)

OpenStudy (anonymous):

the square root grows slower so the limit will be zero

OpenStudy (agent0smith):

Divide both numerator and denominator by x, to solve it algebraically.

OpenStudy (christos):

I am sorry the actual epression is sort(x-3/x)

OpenStudy (christos):

sqrt((x-3)/x)

OpenStudy (anonymous):

\[\lim_{x\to \infty}\sqrt{\frac{x-3}{x}}\]?

OpenStudy (christos):

Yes thats correct

OpenStudy (anonymous):

then it is one since \(x-3\) and \(x\) grow at the same rate

OpenStudy (anonymous):

think what it would be like if \(x\) was \(1,000\) you would have \[\sqrt{.993}\]

OpenStudy (agent0smith):

\[\Large \lim_{x\to \infty}\sqrt{\frac{x-3}{x}} = \sqrt {\lim_{x\to \infty}\frac{x-3} {x}}\]

OpenStudy (christos):

bro I am trying to solve this limit actually http://screencast.com/t/Yoo5IO0bRPjD And I ended up in the expression above, but what you give me isnt the final answer, something must be wrong :(

OpenStudy (agent0smith):

@Christos that's a completely different expression... break it up using limit laws.

OpenStudy (christos):

@satellite73

OpenStudy (christos):

ohh let me try hold on

OpenStudy (christos):

What do you mean by limit laws?

OpenStudy (agent0smith):

\[\Large = \lim_{x \rightarrow \infty} \sqrt{x^2 - 3x} -\lim_{x \rightarrow \infty} x\]

OpenStudy (christos):

so thats infinity - infinity - infinity = +infinity?

OpenStudy (agent0smith):

From a graph it looks like it'll be -1.5, I'm just trying to find a way to show that algebraically.

OpenStudy (christos):

It's ok man dont brother I found it thank you

OpenStudy (agent0smith):

Yeah i'm glad i didn't try to do it myself... wolfram gives a looooong solution: http://www3.wolframalpha.com/Calculate/MSP/MSP35891gdgbb0f2cc1ecie00005387i303c364idgg?MSPStoreType=image/png&s=42&w=433&h=1816